### v1

# Efficient Reasoning via Thought-Training and Thought-Free Inference

Abstract
Recent advances in large language models (LLMs) have leveraged explicit Chain-of-Thought (CoT) prompting to improve reasoning accuracy. However, most existing methods primarily compress verbose reasoning outputs. These Long-to-Short transformations aim to improve efficiency, but still rely on explicit reasoning during inference. In this work, we introduce 3TF (Thought-Training and Thought-Free inference), a framework for efficient reasoning that takes a Short-to-Long perspective. We first train a hybrid model that can operate in both reasoning and non-reasoning modes, and then further train it on CoT-annotated data to internalize structured reasoning, while enforcing concise, thought-free outputs at inference time using the no-reasoning mode. Unlike compression-based approaches, 3TF improves the reasoning quality of non-reasoning outputs, enabling models to perform rich internal reasoning implicitly while keeping external outputs short. Empirically, 3TF-trained models obtain large improvements on reasoning benchmarks under thought-free inference, demonstrating that high quality reasoning can be learned and executed implicitly without explicit step-by-step generation.
Efficient Reasoning via Thought-Training and Thought-Free Inference
Canhui Wu1,2thanks: Work completed during internship at JD.com, Qiong Cao2thanks: Corresponding author, Chao Xue2, Wei Xi1, Xiaodong He2 1Xi’an Jiaotong University,  2JD Future Academy wucanhui@stu.xjtu.edu.cncaoqiong1@jd.com
1 Introduction
Large language models (LLMs) have demonstrated remarkable reasoning abilities across arithmetic, commonsense, and symbolic domains. Chain-of-Thought (CoT) prompting wei2022chain; kojima2022large further enhances these abilities by encouraging models to decompose complex problems into intermediate reasoning steps. This explicit reasoning paradigm has proven crucial for improving model interpretability and success rates on multi-step reasoning tasks.
However, recent studies have revealed a growing inefficiency in the reasoning process of advanced LLMs. When solving simple problems such as “What is 2+3?”, models may produce unnecessarily lengthy explanations or enter cycles of self-reflection—a phenomenon known as overthinking sui2025stop; chen2024not. Overthinking not only increases token consumption and latency aytes2025sketch, but also risks introducing compounding reasoning errors through redundant or self-contradictory steps cuadron2025danger; su2025between. As reasoning-oriented models continue to scale, mitigating this inefficiency has become a critical research focus.
To improve reasoning efficiency, existing approaches generally fall into three categories. Prompt-based methods guide models to produce shorter responses through concise or controlled prompting renze2024benefits; xu2025chain. While simple and effective in some cases, these methods rely heavily on handcrafted instructions and often fail to generalize across tasks. Reinforcement learning (RL)-based strategies offer a more adaptive solution by rewarding correctness while penalizing excessive token usage luo2025o1; yi2025shorterbetter; arora2025traininglanguagemodelsreason; hou2025thinkprune. Supervised fine-tuning (SFT) approaches attempt to train models on datasets with concise reasoning traces, promoting brevity through direct supervision ma2025cot; kang2025c3ot; xia2025tokenskip; yu2025long. All these methods can be viewed as Long-to-Short approaches, aiming to reduce reasoning length while maintaining performance.
In this work, we propose 3TF, a new framework that approaches efficient reasoning from a Short-to-Long perspective. Instead of compressing explicit reasoning traces, 3TF employs an asymmetric training and inference paradigm: during training, the model learns from CoT-annotated data to internalize reasoning patterns; during inference, it operates in a thought-free mode, directly outputting concise final answers. This design enables the model to retain the benefits of CoT supervision—structured internal reasoning—while avoiding explicit, verbose outputs that cause overthinking.
Unlike prior methods that explicitly constrain reasoning length, 3TF focuses on implicit reasoning transfer. By learning from detailed CoT annotations during training, the model internalizes structured reasoning patterns that can be activated silently at inference, transforming reasoning from an explicit linguistic process into a latent cognitive one. This approach allows the model to adaptively allocate reasoning effort: simple tasks can be solved with minimal internal computation, while complex problems trigger richer latent reasoning without producing verbose intermediate steps.
We implement 3TF on both hybrid reasoning models (HRMs) and base models, demonstrating its effectiveness across architectures with different reasoning priors. These models can flexibly operate between explicit and implicit reasoning modes, allowing concise, thought-free inference while preserving the structured reasoning patterns learned from CoT-annotated training data. This unified design enables controlled reasoning efficiency and adaptability to tasks of varying difficulty, and it naturally supports scaling, as larger models can recover more reasoning capability internally without additional supervision.
Extensive experiments show that 3TF consistently maintains high reasoning accuracy while substantially reducing output length. On GSM8K and MATH500, it retains over 90% accuracy compared to full-CoT baselines while using roughly one third of the tokens; on harder benchmarks such as OLYMPIC and AIME24, it preserves 70–80% of full-CoT performance with significantly shorter outputs. These results indicate that effective reasoning can be internalized and executed implicitly, without requiring explicit intermediate steps, highlighting that high-quality reasoning need not always be verbalized to be effective.
Our contributions are summarized as follows:
  1. 1.
    We propose 3TF, an asymmetric Short-to-Long framework that decouples training and inference, enabling LLMs to internalize reasoning via CoT supervision while producing concise, thought-free outputs.
  2. 2.
    We show that reasoning-only training strengthens latent reasoning priors and improves implicit reasoning recoverability as model scale increases.
  3. 3.
    We demonstrate that 3TF achieves state-of-the-art efficiency–accuracy trade-offs across multiple reasoning benchmarks, offering a new paradigm for efficient reasoning in LLMs.
2 Related Work
2.1 Reasoning in LLMs
Large language models (LLMs) have demonstrated remarkable capabilities in solving complex reasoning tasks, particularly when guided by carefully designed prompting strategies. A key advancement in this domain is Chain-of-Thought (CoT) prompting wei2022chain; kojima2022large; wang2022self, which encourages models to generate intermediate reasoning steps before producing a final answer. Subsequent works have shown that CoT improves performance across various reasoning benchmarks. Techniques such as self-consistency decoding wang2022self and multi-step verification creswell2022selection further refine CoT by aggregating multiple reasoning paths, leading to more robust and accurate outputs. To enhance these abilities even further, leading LLMs such as OpenAI-O1 jaech2024openai, DeepSeek-R1 guo2025deepseek, and QwQ team2025qwq integrate advanced methods including reinforcement learning and multi-stage training. These strategies enable detailed multi-step reasoning, achieving state-of-the-art performance in fields such as advanced mathematics and competitive programming. Despite its success, CoT inference introduces substantial computational overhead because the reasoning traces must be explicitly generated at inference time, motivating research into more efficient approaches that retain high reasoning performance.
2.2 Efficient Reasoning
Figure 1: Overview of the 3TF Asymmetric Reasoning Paradigm. (Top) The templates used for training and inference. (Bottom) The training and inference flowchart. A Base model is first trained on a mix of Think and No-Think data to create a Hybrid Model. This model is then further fine-tuned using the Think data. For No-Think Inference, the model is prompted with the Inference Template, which suppresses the explicit reasoning trace and compels the model to generate only the final answer.
As reasoning in LLMs becomes more verbose, recent works aim to make it more concise while preserving quality and accuracy. Prompt-based methods, such as CCoT renze2024benefits, guide models with explicit directives like “Be concise,” while CoD xu2025chain and Token-budget han2024token enforce strict token limits to prevent excessively long outputs. SFT-based approaches, including C3ot kang2025c3ot, Cot-Valve ma2025cot, TokenSkip xia2025tokenskip, and LS-Mixture yu2025long, fine-tune models on datasets containing reasoning of varying lengths, with particular emphasis on shorter and more efficient chains. Merge-based methods wu2025unlocking combine reasoning and instruction model parameters to reduce length efficiently without additional training, while RL-based techniques like DAST shen2025dast, O1-pruner luo2025o1, ShorterBetter yi2025shorterbetter, and TrainEfficient arora2025traininglanguagemodelsreason introduce length penalties during training to encourage concise outputs. Other work focuses on compressing reasoning, for example by distilling system 2 reasoning into system 1 yu2024distilling, preserving reasoning capabilities while significantly reducing inference complexity.
3 From Explicit Thoughts to Implicit Reasoning
3.1 Motivation and Intuition
Recent hybrid reasoning models demonstrate that a single model can flexibly alternate between generating explicit thoughts when necessary and answering directly when possible. This adaptability highlights a growing interest in efficient reasoning that does not always require overt verbalization. However, explicit reasoning often remains verbose, unstable, and computationally expensive.
We propose 3TF (Thought-Training and Thought-Free inference), an asymmetric paradigm that decouples reasoning behavior between training and inference. During training, the model learns from reasoning-augmented data containing both explicit thoughts and final answers; during inference, it is explicitly constrained to suppress reasoning and output only the final answer. Hence, 3TF leverages reasoning supervision during training while enforcing thought-free execution at test time. This asymmetry provides a principled balance between interpretability during learning and efficiency during deployment.
3.2 Asymmetric Reasoning Process
We denote each training instance as a triplet $(x,t,a)$, where $x$ represents the input (e.g., a question or problem description), $t$ denotes the reasoning trace that explicates the intermediate thought process, and $a$ corresponds to the final answer. During training, the model learns to approximate the joint distribution of reasoning and answer generation conditioned on the input:
$P_{\theta}(t,a\mid x)=P_{\theta}(t\mid x)\,P_{\theta}(a\mid t,x),$ (1)
where the first component $P_{\theta}(t\mid x)$ models the generation of reasoning steps, and the second component $P_{\theta}(a\mid t,x)$ models the derivation of the final answer given both the input and the reasoning trace. This decomposition encourages the model to explicitly learn structured reasoning behavior, aligning its internal computation with interpretable chains of thought.
However, the inference stage deliberately adopts an asymmetric process. Unlike training, where explicit reasoning tokens are available, inference suppresses their textual emission. Formally, we denote the inference distribution as conditioning on an empty reasoning context:
$P_{\theta}(a\mid x,t=\varnothing),$ (2)
where $t=\varnothing$ is not a random variable from the training distribution but an abstract symbol representing the fixed, empty reasoning context at inference. Although $P_{\theta}(a\mid x,t=\varnothing)$ is not directly optimized during training, the model’s behavior under this condition reflects its capacity for implicit reasoning.
During deployment, decoding is initialized with an empty reasoning field such that the model is conditioned on $t=\varnothing$ from the outset. This simple structural constraint ensures that the model performs reasoning implicitly while emitting only the final answer, maintaining internal coherence without exposing intermediate steps.
Such an asymmetric design achieves two essential goals: (1) it retains the reasoning competence acquired through chain-of-thought supervision, and (2) it suppresses redundant or unstable intermediate outputs, improving inference efficiency and stability. Thus, 3TF provides controllable, interpretable, and computationally efficient reasoning behavior without requiring explicit traces at test time.
3.3 Information-Theoretic Explanation
From an information-theoretic perspective, explicit and implicit reasoning correspond to two complementary modes of information transmission between the input $x$ and the final answer $a$: expansion through intermediate reasoning traces and compression through direct answering.
Reasoning as Expansion.
Explicit reasoning unfolds the information contained in $x$ into a higher-bandwidth intermediate representation $t$, forming an information chain $x\rightarrow t\rightarrow a$. Since $t$ captures fine-grained dependencies and latent structure, it generally holds that $I(x;t)\gg I(x;a)$. Thus, reasoning acts as an expansion operator: it temporarily increases the representational capacity so that the subsequent generation of $a$ becomes more deterministic and structured.
Training as Alignment.
During training, the model optimizes the joint process in Equation (1) to maximize the mutual information between input, reasoning, and answer. This can be viewed as maximizing
$\max_{\theta}\;I_{\theta}(x;t)+I_{\theta}(t;a\mid x),$ (3)
which encourages $t$ to preserve input information while being maximally predictive of $a$. Reasoning therefore serves as a structured channel through which the model learns to reorganize and align information before final compression.
Inference as Controlled Compression.
At inference, when reasoning emission is suppressed $(t=\varnothing)$, the model effectively compresses the same latent structure directly into $a$. Since training has already aligned reasoning and answering channels, this compression can maintain high fidelity even in the absence of explicit reasoning tokens. Hence, implicit reasoning operates as an information bottleneck that preserves essential semantics while minimizing representational overhead.
Adaptive Behavior.
Finally, because the information-control mechanism is internally learned, the model dynamically adjusts its implicit reasoning depth according to task complexity: Simple queries lead to minimal compression (short answers), while complex problems elicit richer internal representations even without external traces. In this sense, 3TF realizes a self-regulating, task-conditioned compression policy that unifies interpretability, efficiency, and adaptability in neural reasoning.
4 Experimental
4.1 Experimental Setup
Models and Datasets
We conduct experiments on multiple sizes of the Qwen3 yang2025qwen3 model, including 4B, 8B, and 14B variants. Qwen3 is a hybrid reasoning model that is well-suited for experimental studies. For training, we randomly sample 100K problems from the MATH subset of AM-Qwen3-Distilled tian2025correctanswersequaldistillation. We evaluate 3TF on four reasoning benchmarks: AIME24, MATH500 lightman2023let, GSM8K cobbe2021gsm8k, and Olympiad.
Methods GSM8K@1 MATH-500@4 OLYMPIC@4 AIME 2024@16
Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio Acc. Tok. Ratio
Qwen3-4B
Thinking 93.7 1792 100% 97.2 4229 100% 80.7 10122 100% 80.0 13200 100%
NoThinking 88.1 250 13.9% 91.6 972 23.0% 63.5 1778 17.6% 43.3 3792 28.7%
BeConcise 93.4 794 44.3% 97.8 3513 83.1% 80.7 8790 86.8% 80.0 13053 98.9%
ChainofDraft 94.3 567 31.6% 96.6 2457 58.1% 80.8 7339 72.5% 83.3 10626 80.5%
CoT-Valve 85.9 314 17.5% 91.4 1106 26.1% 61.8 1644 16.2% 53.3 1870 14.2%
LS-Mixture 93.1 1702 95.0% 96.2 3672 86.8% 78.5 8772 86.6% 73.3 12670 95.9%
3TF 92.4 278 15.5% 95.6 1505 35.6% 73.0 4097 40.5% 70.0 9123 69.1%
Qwen3-8B
Thinking 94.2 2020 100% 98.0 4894 100% 81.4 10454 100% 80.0 13173 100%
NoThinking 89.9 245 12.1% 93.2 1096 22.4% 65.5 2081 19.9% 50.0 3581 27.2%
BeConcise 95.0 946 46.8% 98.8 4010 82.0% 81.6 9695 92.7% 83.3 12772 97.0%
ChainofDraft 94.0 500 24.8% 97.2 2978 60.8% 81.7 8386 80.2% 83.3 11365 86.3%
CoT-Valve 88.1 318 15.7% 93.2 764 15.6% 65.3 1934 18.5% 53.3 1916 14.5%
LS-Mixture 93.7 1563 77.4% 96.9 3735 76.3% 80.3 8382 80.2% 80.0 13091 99.4%
3TF 93.0 241 11.9% 96.6 1358 27.7% 76.0 3134 30.0% 76.6 7961 60.4%
Qwen3-14B
Thinking 94.6 1525 100% 98.8 3748 100% 83.9 9675 100% 90.0 13253 100%
NoThinking 92.1 247 16.2% 94.4 803 21.4% 67.3 1791 18.5% 56.6 3153 23.8%
BeConcise 94.4 656 43.0% 99.0 3206 85.6% 83.0 8511 88.0% 90.0 12777 96.4%
ChainofDraft 95.3 585 38.4% 97.0 2286 61.0% 82.0 7006 72.4% 86.6 10441 78.7%
CoT-Valve 90.0 311 20.4% 94.0 814 21.7% 65.7 1510 15.6% 50.0 2105 15.9%
LS-Mixture 93.7 1248 81.8% 97.4 3281 87.5% 78.4 6788 70.1% 83.3 10584 79.8%
3TF 94.4 275 18.0% 98.6 1613 43.0% 79.4 4968 51.3% 86.6 10280 77.5%
Table 1: Experimental results of 3TF on reasoning models of various sizes. For CoT-Valve and LS-mixture, we fine-tune the models on their official datasets.
Implementation Details
We fine-tune the Qwen3 model using the Swift zhao2024swiftascalablelightweightinfrastructure framework for 3 epochs with a batch size of 32. The learning rate is set to 2e-5 with a warmup ratio of 0.05. We enable sequence packing to improve training efficiency. We use lighteval lighteval for evaluation. The number of samplings during evaluation depends on the dataset: 16 samples per question for AIME 2024, 4 samples per question for MATH500 and Olympiad, and 1 sample per question for GSM8K.
Baselines
We compare our method against three types of baselines. 1) Qwen3 Modes: The original Qwen3 model is evaluated under both the Think and No-Think modes to measure the impact of explicit reasoning instructions. 2) Prompt-based Responses: We evaluate prompts that encourage the model to produce concise answers, including the BeConcise renze2024benefits and ChainofDraft xu2025chain strategies, which aim to streamline output while retaining correctness. 3) SFT with Compressed Reasoning Data: We include baselines trained via supervised fine-tuning on datasets augmented with compressed reasoning chains, such as CoT-Valve ma2025cot and LS-Mixture yu2025long, to assess the benefit of incorporating distilled reasoning steps.
4.2 Comparison with Baselines
Figure 2: Accuracy scaling of 3TF from 4B to 32B across four reasoning benchmarks. As model size increases, 3TF progressively closes the gap to the Base-Think mode, showing improved reasoning recoverability at scale.
Figure 3: Avg@k performance scaling of 3TF from 4B to 32B across four reasoning benchmarks. Similar to the pass@k accuracy results, 3TF progressively closes the gap to the Base-Think mode, showing that reasoning recoverability for this metric also improves with scale.
Figure 4: Impact of mixing No-Think data during training on Qwen3-8B across four reasoning benchmarks. Increasing the No-Think ratio consistently shortens model outputs (orange) but monotonically reduces accuracy (blue). Even 1% No-Think materially hurts reasoning fidelity, indicating high sensitivity to non-reasoning supervision.
Table 1 reports the results of 3TF and all baselines across four reasoning benchmarks and three model scales. Overall, 3TF achieves a strong trade-off between reasoning accuracy and output efficiency.
Overall Performance.
Across all datasets, 3TF attains accuracy comparable to the full Thinking mode while producing substantially shorter responses. For instance, on Qwen3-14B, 3TF achieves 94.4% on GSM8K and 97.8% on MATH500, nearly matching the full-CoT performance (94.6% and 97.8%) but using only 19% and 43% of the output tokens on GSM8K and MATH500, respectively. Similar trends hold across smaller model variants, showing that the proposed asymmetric training–inference design effectively preserves reasoning ability without explicit reasoning traces. Moreover, across datasets of varying difficulty, 3TF demonstrates strong adaptability—maintaining concise outputs on GSM8K while exhibiting a lower compression ratio on more challenging benchmarks such as AIME.
Comparison.
Prompt-based methods generally maintain the original model performance in terms of accuracy, but achieve lower compression ratios compared to 3TF. This gap is particularly evident on simpler datasets such as GSM8K and MATH500, where 3TF produces much more compact outputs. For fine-tuning based methods, CoT-Valve achieves high compression but suffers a significant drop in performance, whereas the model trained with LM-Mixture maintains accuracy but yields a much lower compression ratio. Taken together, 3TF achieves a strong trade-off between accuracy and output compression.
4.3 Scaling Law of Reasoning Recoverability
As shown in Figure 2, we analyze the scaling behavior of 3TF across model sizes from 4B to 32B parameters. A clear monotonic trend emerges: as the parameter count increases, the performance gap between 3TF and the base_think mode consistently narrows across all four benchmarks. On GSM8K and MATH500, the trained No-Think model at 14B already nearly saturates full CoT performance, and at 32B the remaining delta is minimal. Even on more challenging distributions such as Olympic and AIME24, where smaller models still show non-trivial degradation, the 32B checkpoint closes most of the gap.
These results highlight an important scaling dynamic. Larger reasoning-centric LMs contain stronger latent reasoning priors that can be efficiently unlocked through asymmetric 3TF training. In other words, as scale increases, reasoning ability becomes more recoverable without explicit chain-of-thought. The implication is not simply that scaling improves raw accuracy, but that scaling structurally increases reasoning accessibility. As a result, the efficiency advantage of 3TF compounds with scale: larger models reason better, while also requiring far less explicit reasoning at inference time to maintain that level of quality.
4.4 Avg@k Results
As shown in Figure 3, 3TF exhibits a consistent scaling trend under Avg@k. Across all four benchmarks, performance improves monotonically with model size, and the gap to Base-Think narrows steadily. On GSM8K and MATH500, 3TF almost matches full CoT at 32B, with residual differences below 1–2%. This suggests that large models can recover most of the reasoning benefit even when inference is executed under no-think.
For harder settings (OLYMPIC, AIME24), smaller models show large degradation relative to Base-Think (up to 30–37% at 4B). Yet scale again closes this gap: at 32B, 3TF recovers more than 10% on OLYMPIC and more than 25% on AIME24. This indicates strong scale-driven reasoning recoverability.
Overall, Avg@k reinforces the same pattern as earlier sections: scale makes reasoning increasingly accessible without explicit chain-of-thought at test time. Practically, 3TF becomes more efficient with scale, paying far fewer tokens for nearly the same effective reasoning competence, while retaining room for sparse targeted reasoning only when required.
4.5 Effect of Adding No-Think Data
We further evaluate the impact of training with No-Think data on Qwen3-8B. Adding No-Think samples consistently shortens outputs across all datasets (e.g., GSM8K average length drops from 241 to 207 tokens, AIME24 from 7961 to 1642). However, this compression is coupled with substantial accuracy loss: GSM8K drops from 93.0% to 90.3%, and AIME24 from 76.6% to 43.3%. MATH500 and OLYMPIC exhibit the same trend. Figure 4 visualizes the effect of different No-Think ratios.
Even extremely small amounts (1% of training) cause noticeable degradation across all benchmarks. Thus, while No-Think data reliably enforces conciseness, excessive or even minimal inclusion harms reasoning fidelity, highlighting the need for careful balancing of reasoning vs non-reasoning samples during training.
4.6 Train from Scratch
Methods GSM8K@1 MATH-500@4 OLYMPIC@4 AIME 2024@16
Acc. Tok. Acc. Tok. Acc. Tok. Acc. Tok.
Qwen2.5-7B-Instruct
NoThinking 80.8 278 86.6 576 50.0 935 26.6 940
Hybrid (Think) 92.6 1834 94.0 7352 67.2 13241 56.6 20331
Hybrid (NoThink) 88.4 310 90.0 1727 56.5 2253 46.6 5023
3TF 91.0 589 94.6 4143 63.0 5260 56.6 12025
Qwen2.5-7B-R1-Distill
Think 86.4 541 96.4 2971 77.6 7690 76.6 10414
Hybrid (Think) 87.0 322 96.0 3601 68.2 2410 76.6 11786
Hybrid (NoThink) 86.8 271 90.0 778 61.8 1200 43.3 2654
3TF 88.6 287 93.6 913 64.2 2008 60.0 4346
Table 2: Results of 3TF applied to Qwen2.5-7B models first trained on mixed Think/NoThink data to form hybrid reasoning, then fine-tuned with reasoning-only samples.
GSM8K MATH Olympic AIME
(1319) (500) (674) (30)
Total Count
Think 61,936 42,722 140,563 6,903
NoThink 485 959 2,134 137
3TF 1,822 1,746 13,120 1,032
Average Count per Question
Think 46.96 85.44 208.55 230.10
NoThink 0.37 1.92 3.17 4.57
3TF 1.38 3.49 19.47 34.40
Table 3: Total count and average count per question of reasoning tokens across four benchmarks under Think, NoThink, and 3TF. Numbers in parentheses indicate the number of questions in each benchmark.
Unlike previous experiments where 3TF was directly fine-tuned on hybrid reasoning models such as Qwen3, here we first train a base model into a hybrid reasoning form using mixed Think and NoThink data, and then apply 3TF training purely on reasoning samples. This setup examines whether asymmetric supervision remains effective without relying on pre-existing hybrid priors.
As shown in Table 2, 3TF consistently achieves high accuracy with concise outputs. On Qwen2.5-7B-Instruct, it reaches 91.0% on GSM8K and 94.6% on MATH500—comparable to the Hybrid (Think) model but using only about one third of its tokens. Similar gains are observed on Qwen2.5-7B-R1-Distill, where 3TF attains 88.6% and 93.6% accuracy on GSM8K and MATH500 while keeping responses compact. The distilled variant generally yields stronger reasoning recoverability; for example, on AIME24, it achieves 60.0% accuracy with less than half the token length of the Instruct version. Overall, 3TF proves effective even when trained from scratch, showing that asymmetric reasoning fine-tuning can independently induce concise yet faithful reasoning behavior.
4.7 Analysis of Output Structure
To understand how 3TF alters the internal structure of model outputs, we compare the frequency of canonical “reasoning words” across Think, No-Think, and 3TF. The lexical indicators used as reasoning proxies (e.g., “wait”, “however”, “check”) are listed in the Appendix. Table 3 reports both total counts and per-question frequencies of these words over four reasoning benchmarks.
3TF increases the use of reasoning markers relative to No-Think, but still remains far below full Think mode, often more than an order of magnitude smaller. Thus 3TF does not revert into implicit full chain-of-thought. Instead, it preserves short, targeted inference traces while retaining most of the efficiency gains.
We also observe that as benchmark difficulty and distribution shift increase (e.g., Olympic, AIME), all modes (including 3TF) emit more reasoning markers and produce longer outputs. Models appear to adaptively expose more minimal reasoning only when demanded by harder tasks. Taken together, 3TF induces a sparse reasoning regime: not pure No-Think, but substantially shorter than Think, surfacing only critical reasoning artifacts when needed.
5 Conclusion
In this work, we presented 3TF, a novel framework for efficient reasoning in large language models. By first training a hybrid model and then further training on chain-of-thought annotated data while enforcing concise, thought-free outputs at inference in the no-reasoning mode, 3TF enables models to internalize rich reasoning patterns while avoiding verbose explanations. Extensive experiments across multiple benchmarks and model scales demonstrate that 3TF preserves reasoning accuracy, reduces output length, and adapts to problem difficulty. Our results highlight a practical approach for scalable and efficient implicit reasoning, offering a new direction for deploying reasoning-capable LLMs with minimal inference overhead.
6 Limitations
While 3TF demonstrates strong performance and efficiency across multiple benchmarks and model scales, there are several limitations worth noting. First, on smaller-scale models (e.g., Qwen3-4B), the pass@1 accuracy remains notably lower compared to larger models, indicating that reasoning recoverability under minimal output sampling is still limited at small scale. Second, although 3TF reduces the number of generated tokens, it may still produce slightly longer outputs on particularly challenging benchmarks (e.g., AIME24 and Olympiad) to preserve accuracy, which partially offsets efficiency gains. Finally, 3TF has been evaluated primarily on mathematics-focused benchmarks, and its generalizability to other domains remains to be thoroughly investigated. Addressing these limitations is an important direction for future work, especially for applying 3TF to smaller models or broader domains.
Appendix A Appendix
A.1 Experimental environment
We trained on the NVIDIA 8*H200 node, with CUDA version 12.7 and CPU model Intel(R) Xeon(R) Platinum 8468V.
A.2 Reasoning Words
The table below lists the reasoning words we use:
Reasoning Words
“wait”, “alternatively”, “but”, “however”,
“alternative”, “check”, “double-check”,
“hmm”, “okay”, “maybe”
A.3 Prompt
The prompt used in the Prompt method is shown below.
CoD Think step by step, but only keep a minimum draft for each thinking step, with 5 words at most.
CCoT be concise

### v3

@@
 Recent advances in large language models (LLMs) have leveraged explicit Chain-of-Thought (CoT) prompting to improve reasoning accuracy.
-However, most existing methods primarily compress verbose reasoning outputs.
-These Long-to-Short transformations aim to improve efficiency, but still rely on explicit reasoning during inference.
+However, most existing methods primarily focus on compressing verbose reasoning outputs.
+These Long-to-Short transformations aim to improve efficiency, but require a large amount of short CoT data.
 In this work, we introduce 3TF (Thought-Training and Thought-Free inference), a framework for efficient reasoning that takes a Short-to-Long perspective.
@@
 In this work, we propose 3TF, a new framework that approaches efficient reasoning from a Short-to-Long perspective.
-Instead of compressing explicit reasoning traces, 3TF employs an asymmetric training and inference paradigm: during training, the model learns from CoT-annotated data to internalize reasoning patterns; during inference, it operates in a thought-free mode, directly outputting concise final answers.
-This design enables the model to retain the benefits of CoT supervision—structured internal reasoning—while avoiding explicit, verbose outputs that cause overthinking.
-Unlike prior methods that explicitly constrain reasoning length, 3TF focuses on implicit reasoning transfer.
-By learning from detailed CoT annotations during training, the model internalizes structured reasoning patterns that can be activated silently at inference, transforming reasoning from an explicit linguistic process into a latent cognitive one.
-This approach allows the model to adaptively allocate reasoning effort: simple tasks can be solved with minimal internal computation, while complex problems trigger richer latent reasoning without producing verbose intermediate steps.
-We implement 3TF on both hybrid reasoning models (HRMs) and base models, demonstrating its effectiveness across architectures with different reasoning priors.
-These models can flexibly operate between explicit and implicit reasoning modes, allowing concise, thought-free inference while preserving the structured reasoning patterns learned from CoT-annotated training data.
-This unified design enables controlled reasoning efficiency and adaptability to tasks of varying difficulty, and it naturally supports scaling, as larger models can recover more reasoning capability internally without additional supervision.
+Instead of compressing explicit reasoning traces, 3TF employs an asymmetric training and inference paradigm.
+During training, the model learns from CoT-annotated data to internalize reasoning patterns.
+During inference, it operates in a thought-free mode, directly outputting concise final answers.
+This design allows the model to retain the benefits of CoT supervision, such as structured internal reasoning, while avoiding explicit and verbose outputs that lead to overthinking.
+Previous SFT approaches typically require access to shortened CoT data for model training, such as summarization via LLMs, manual rewriting, or model merging, which can be costly and time-consuming.
+Unlike prior methods that explicitly constrain reasoning length, 3TF emphasizes implicit reasoning transfer.
+The training of 3TF is conducted in two stages: in the first stage, we train a hybrid reasoning model (HRM) capable of flexibly switching between explicit and implicit reasoning modes.
+In the second stage, the model learns exclusively from detailed CoT data to further strengthen its reasoning capability.
+The resulting model internalizes structured reasoning patterns, which can be silently activated during inference, effectively transforming reasoning from an explicit linguistic process into a latent cognitive one.
 Extensive experiments show that 3TF consistently maintains high reasoning accuracy while substantially reducing output length.
@@
 Figure 1: Overview of the 3TF Asymmetric Reasoning Paradigm.
-(Top) The templates used for training and inference.
-(Bottom) The training and inference flowchart.
-A Base model is first trained on a mix of Think and No-Think data to create a Hybrid Model.
-This model is then further fine-tuned using the Think data.
-For No-Think Inference, the model is prompted with the Inference Template, which suppresses the explicit reasoning trace and compels the model to generate only the final answer.
+(Top) Training and inference templates.
+(Bottom) Flowchart: A Base model is trained on both Think and No-Think data, then fine-tuned on Think data.
+For No-Think Inference, the model uses the Inference Template to generate only the final answer.
 As reasoning in LLMs becomes more verbose, recent works aim to make it more concise while preserving quality and accuracy.
@@
 3.1 Motivation and Intuition
-Recent hybrid reasoning models demonstrate that a single model can flexibly alternate between generating explicit thoughts when necessary and answering directly when possible.
-This adaptability highlights a growing interest in efficient reasoning that does not always require overt verbalization.
+Recent hybrid reasoning models demonstrate the ability to alternate flexibly between explicit reasoning and direct answers, highlighting a growing interest in efficient reasoning that does not always require overt verbalization.
 However, explicit reasoning often remains verbose, unstable, and computationally expensive.
-We propose 3TF (Thought-Training and Thought-Free inference), an asymmetric paradigm that decouples reasoning behavior between training and inference.
-During training, the model learns from reasoning-augmented data containing both explicit thoughts and final answers; during inference, it is explicitly constrained to suppress reasoning and output only the final answer.
-Hence, 3TF leverages reasoning supervision during training while enforcing thought-free execution at test time.
-This asymmetry provides a principled balance between interpretability during learning and efficiency during deployment.
+To address this, we propose 3TF (Thought-Training and Thought-Free inference), an asymmetric paradigm that decouples the reasoning process during training and inference.
+During training, the model learns from reasoning-augmented data, which includes both explicit reasoning steps and final answers, while at inference time, it is explicitly constrained to suppress reasoning and output only the final answer.
+This approach enables the model to benefit from reasoning supervision during training, while enforcing thought-free execution during deployment.
 3.2 Asymmetric Reasoning Process
@@
 3.3 Information-Theoretic Explanation
-From an information-theoretic perspective, explicit and implicit reasoning correspond to two complementary modes of information transmission between the input $x$ and the final answer $a$: expansion through intermediate reasoning traces and compression through direct answering.
-Reasoning as Expansion.
-Explicit reasoning unfolds the information contained in $x$ into a higher-bandwidth intermediate representation $t$, forming an information chain $x\rightarrow t\rightarrow a$.
-Since $t$ captures fine-grained dependencies and latent structure, it generally holds that $I(x;t)\gg I(x;a)$.
-Thus, reasoning acts as an expansion operator: it temporarily increases the representational capacity so that the subsequent generation of $a$ becomes more deterministic and structured.
-Training as Alignment.
-During training, the model optimizes the joint process in Equation (1) to maximize the mutual information between input, reasoning, and answer. This can be viewed as maximizing
-$\max_{\theta}\;I_{\theta}(x;t)+I_{\theta}(t;a\mid x),$ (3)
-which encourages $t$ to preserve input information while being maximally predictive of $a$.
-Reasoning therefore serves as a structured channel through which the model learns to reorganize and align information before final compression.
-Inference as Controlled Compression.
-At inference, when reasoning emission is suppressed $(t=\varnothing)$, the model effectively compresses the same latent structure directly into $a$.
-Since training has already aligned reasoning and answering channels, this compression can maintain high fidelity even in the absence of explicit reasoning tokens.
-Hence, implicit reasoning operates as an information bottleneck that preserves essential semantics while minimizing representational overhead.
-Adaptive Behavior.
-Finally, because the information-control mechanism is internally learned, the model dynamically adjusts its implicit reasoning depth according to task complexity: Simple queries lead to minimal compression (short answers), while complex problems elicit richer internal representations even without external traces.
-In this sense, 3TF realizes a self-regulating, task-conditioned compression policy that unifies interpretability, efficiency, and adaptability in neural reasoning.
+From an information-theoretic view, explicit and implicit reasoning are complementary.
+Explicit reasoning expands input $x$ into intermediate traces $t$ ($x\rightarrow t\rightarrow a$), capturing fine-grained dependencies ($I(x;t)\gg I(x;a)$), while training aligns $t$ with $a$ via
+$\max_{\theta}I_{\theta}(x;t)+I_{\theta}(t;a\mid x).$
+At inference, suppressing $t$ compresses information directly into $a$, leveraging the aligned channels to retain essential semantics.
+Explicit reasoning expands and structures information; implicit reasoning compresses it efficiently.
 4 Experimental
@@
 The number of samplings during evaluation depends on the dataset: 16 samples per question for AIME 2024, 4 samples per question for MATH500 and Olympiad, and 1 sample per question for GSM8K.
+We compute pass@k based on these sampled generations.
 Baselines
@@
 Across all datasets, 3TF attains accuracy comparable to the full Thinking mode while producing substantially shorter responses.
-For instance, on Qwen3-14B, 3TF achieves 94.4% on GSM8K and 97.8% on MATH500, nearly matching the full-CoT performance (94.6% and 97.8%) but using only 19% and 43% of the output tokens on GSM8K and MATH500, respectively.
+For instance, on Qwen3-14B, 3TF achieves 94.4% on GSM8K and 97.8% on MATH500, nearly matching the full-CoT performance but using only 19% and 43% of the output tokens on GSM8K and MATH500, respectively.
 Similar trends hold across smaller model variants, showing that the proposed asymmetric training–inference design effectively preserves reasoning ability without explicit reasoning traces.
-Moreover, across datasets of varying difficulty, 3TF demonstrates strong adaptability—maintaining concise outputs on GSM8K while exhibiting a lower compression ratio on more challenging benchmarks such as AIME.
+3TF further adapts its output length to task complexity, remaining highly concise on easier datasets while allowing moderately longer responses on more challenging ones.
 Comparison.
@@
 Taken together, 3TF achieves a strong trade-off between accuracy and output compression.
+Takeaway #1 Training on reasoning-rich data alone can markedly improve the performance of non-thinking inference modes while keeping outputs far more concise than full reasoning traces.
 4.3 Scaling Law of Reasoning Recoverability
@@
 Even on more challenging distributions such as Olympic and AIME24, where smaller models still show non-trivial degradation, the 32B checkpoint closes most of the gap.
-These results highlight an important scaling dynamic.
-Larger reasoning-centric LMs contain stronger latent reasoning priors that can be efficiently unlocked through asymmetric 3TF training.
-In other words, as scale increases, reasoning ability becomes more recoverable without explicit chain-of-thought.
-The implication is not simply that scaling improves raw accuracy, but that scaling structurally increases reasoning accessibility.
-As a result, the efficiency advantage of 3TF compounds with scale: larger models reason better, while also requiring far less explicit reasoning at inference time to maintain that level of quality.
+These results highlight a key scaling dynamic: larger reasoning-centric LMs have stronger latent reasoning priors that can be efficiently unlocked via asymmetric 3TF training.
+As scale increases, reasoning ability becomes more accessible without explicit chain-of-thought.
+Takeaway #2 Scaling up the model steadily narrows the gap between 3TF and full Thinking mode, bringing larger models closer to full reasoning performance.
 4.4 Avg@k Results
-As shown in Figure 3, 3TF exhibits a consistent scaling trend under Avg@k.
-Across all four benchmarks, performance improves monotonically with model size, and the gap to Base-Think narrows steadily.
-On GSM8K and MATH500, 3TF almost matches full CoT at 32B, with residual differences below 1–2%.
-This suggests that large models can recover most of the reasoning benefit even when inference is executed under no-think.
-For harder settings (OLYMPIC, AIME24), smaller models show large degradation relative to Base-Think (up to 30–37% at 4B).
-Yet scale again closes this gap: at 32B, 3TF recovers more than 10% on OLYMPIC and more than 25% on AIME24.
-This indicates strong scale-driven reasoning recoverability.
-Overall, Avg@k reinforces the same pattern as earlier sections: scale makes reasoning increasingly accessible without explicit chain-of-thought at test time.
-Practically, 3TF becomes more efficient with scale, paying far fewer tokens for nearly the same effective reasoning competence, while retaining room for sparse targeted reasoning only when required.
+As shown in Figure 3, 3TF demonstrates a consistent scaling trend under Avg@k.
+Across all four benchmarks, performance improves monotonically with model size, narrowing the gap to Base-Think.
+On GSM8K and MATH500, 3TF at 32B nearly matches full CoT, with differences under 1–2%, suggesting that larger models recover most of the reasoning benefits even without explicit chain-of-thought during inference.
+In more challenging settings (OLYMPIC, AIME24), smaller models show significant performance degradation compared to Base-Think (up to 30–37% at 4B).
+However, scaling mitigates this gap: at 32B, 3TF recovers over 10% on OLYMPIC and more than 25% on AIME24, highlighting the strong recoverability of reasoning through scale.
+Takeaway #3 For smaller models, 3TF under Avg@k lags behind the Thinking mode, but this gap steadily diminishes as model scale increases.
 4.5 Effect of Adding No-Think Data
 We further evaluate the impact of training with No-Think data on Qwen3-8B.
-Adding No-Think samples consistently shortens outputs across all datasets (e.g., GSM8K average length drops from 241 to 207 tokens, AIME24 from 7961 to 1642).
+Adding No-Think samples consistently shortens outputs across all datasets (e.g., AIME24 from 7961 to 1642).
 However, this compression is coupled with substantial accuracy loss: GSM8K drops from 93.0% to 90.3%, and AIME24 from 76.6% to 43.3%.
@@
 Even extremely small amounts (1% of training) cause noticeable degradation across all benchmarks.
-Thus, while No-Think data reliably enforces conciseness, excessive or even minimal inclusion harms reasoning fidelity, highlighting the need for careful balancing of reasoning vs non-reasoning samples during training.
+Thus, while No-Think data reliably enforces conciseness, we hypothesize that the inclusion of non-thought data causes the model to revert to a purely non-reasoning mode, which undermines its ability to perform complex reasoning tasks and leads to performance degradation.
+Takeaway #4 Adding No-Think data significantly reduces 3TF performance.
+4.6 Ablation Study on the Datasets
+Methods GSM8K@1 MATH-500@4
+Acc. Tok. Acc. Tok.
+Qwen3-8B
+NoThink 89.9 245 93.2 1092
+NoThink-T1 90.0 300 91.2 741
+NoThink-T2 90.2 250 93.4 1038
+3TF 93.0 241 96.6 1358
+Table 2: Ablation study on Qwen3-8B comparing different training methods: NoThink, NoThink-T1, NoThink-T2, and 3TF.
+In this section, we compare 3TF with two training strategies: NoThink-T1, which uses thought data but excludes reasoning part, and NoThink-T2, which is trained from distilled non-thought data (from Qwen3 Non-Think mode).
+While NoThink-T1 results in slightly shorter outputs, its performance suffers noticeably compared to 3TF.
+On the other hand, NoThink-T2, trained directly from non-thought data, shows performance comparable to the baseline non-thought model, with no significant improvements.
+These results suggest that 3TF effectively combines reasoning supervision with thought-free inference, outperforming methods that prioritize direct fine-tuning with non-thought data.
+Takeaway #5 Data containing reasoning improves the performance of No-Think modes, rather than the answer content alone.
 4.7 Train from Scratch
@@
 Table 4: Total count and average count per question of reasoning tokens across four benchmarks under Think, NoThink, and 3TF. Numbers in parentheses indicate the number of questions in each benchmark.
-Unlike previous experiments where 3TF was directly fine-tuned on hybrid reasoning models such as Qwen3, here we first train a base model into a hybrid reasoning form using mixed Think and NoThink data, and then apply 3TF training purely on reasoning samples.
-This setup examines whether asymmetric supervision remains effective without relying on pre-existing hybrid priors.
+Unlike previous experiments where 3TF was fine-tuned directly on hybrid reasoning models like Qwen3, we first train a base model into a hybrid reasoning form using mixed Think and NoThink data.
+We then apply 3TF training solely on reasoning samples.
+This setup tests whether asymmetric supervision can be effective without relying on pre-existing hybrid priors.
 As shown in Table 3, 3TF consistently achieves high accuracy with concise outputs.
-On Qwen2.5-7B-Instruct, it reaches 91.0% on GSM8K and 94.6% on MATH500—comparable to the Hybrid (Think) model but using only about one third of its tokens.
-Similar gains are observed on Qwen2.5-7B-R1-Distill, where 3TF attains 88.6% and 93.6% accuracy on GSM8K and MATH500 while keeping responses compact.
-The distilled variant generally yields stronger reasoning recoverability; for example, on AIME24, it achieves 60.0% accuracy with less than half the token length of the Instruct version.
-Overall, 3TF proves effective even when trained from scratch, showing that asymmetric reasoning fine-tuning can independently induce concise yet faithful reasoning behavior.
+On Qwen2.5-7B-Instruct, it reaches 91.0% on GSM8K and 94.6% on MATH500, comparable to the Hybrid (Think) model while using only about a third of its tokens.
+Similar improvements are observed on Qwen2.5-7B-R1-Distill, where 3TF achieves 88.6% and 93.6% accuracy on GSM8K and MATH500 with compact responses.
+The distilled variant shows stronger reasoning recoverability; on AIME24, it reaches 60.0% accuracy with less than half the token length of the Instruct version.
+Overall, 3TF proves effective even when trained from scratch, demonstrating that asymmetric reasoning fine-tuning can independently induce concise yet faithful reasoning behavior.
 4.8 Analysis of Output Structure
-To understand how 3TF alters the internal structure of model outputs, we compare the frequency of canonical “reasoning words” across Think, No-Think, and 3TF.
-The lexical indicators used as reasoning proxies (e.g., “wait”, “however”, “check”) are listed in the Appendix.
-Table 3 reports both total counts and per-question frequencies of these words over four reasoning benchmarks.
-3TF increases the use of reasoning markers relative to No-Think, but still remains far below full Think mode, often more than an order of magnitude smaller.
-Thus 3TF does not revert into implicit full chain-of-thought.
-Instead, it preserves short, targeted inference traces while retaining most of the efficiency gains.
-We also observe that as benchmark difficulty and distribution shift increase (e.g., Olympic, AIME), all modes (including 3TF) emit more reasoning markers and produce longer outputs.
-Models appear to adaptively expose more minimal reasoning only when demanded by harder tasks.
-Taken together, 3TF induces a sparse reasoning regime: not pure No-Think, but substantially shorter than Think, surfacing only critical reasoning artifacts when needed.
+To understand how 3TF alters the internal structure of model outputs, we compare the frequency of canonical "reasoning words" across Think, No-Think, and 3TF.
+These reasoning proxies (e.g., "wait", "however", "check") are listed in the Appendix, with Table 4 reporting both total counts and per-question frequencies across four reasoning benchmarks.
+3TF increases the use of reasoning markers relative to No-Think but remains far below Think, often by more than an order of magnitude.
+This suggests that 3TF does not fully revert to chain-of-thought reasoning, instead maintaining concise, targeted inference traces for efficiency.
+As benchmark difficulty increases (e.g., Olympic, AIME), all modes, including 3TF, emit more reasoning markers and generate longer outputs, reflecting adaptive reasoning exposure based on task complexity.
+Thus, 3TF induces a sparse reasoning regime: it is not pure No-Think but is significantly shorter than Think, only surfacing critical reasoning when needed.
+Takeaway #6 3TF introduces limited reasoning into non-thinking mode and adapts: the harder the task, the more reasoning patterns are expressed.
 5 Conclusion